# Fewer supermassive binary black holes in pulsar timing array observations

### v2

@@ Abstract
We find the inferred properties of the putative gravitational wave background in the second data release of the European Pulsar Timing Array (EPTA) to be in better agreement with theoretical expectations under the improved noise model. In particular, if the gravitational wave background is from supermassive black hole binaries, the observed gravitational wave emission is the dominant source of the binary energy loss, with no evidence of environmental effects, eccentric orbits, or an overabundance of sources. Our improved noise model addresses the misspecification of noise priors which impacts the inference of temporal correlations of the background. We also find evidence for epoch-correlated temporally-uncorrelated (white) noise which has a stronger impact on the measurement in the presence of Hellings-Downs correlations of the background.
@@
Since 2020, Pulsar Timing Arrays (PTAs) have reported growing evidence for the nanohertz-frequency gravitational wave background in their data. The first tentative evidence came from a temporally-correlated stochastic process in pulsar timing data Nanograv Collaboration (2020); Goncharov et al. (2021); Chen et al. (2021); Goncharov et al. (2022). The Fourier spectrum of delays and advances in pulsar pulse arrival times exhibited the expected spectral properties of the background. Most recently, PTAs — with varying levels of statistical significance Nanograv Collaboration (2023a); EPTA Collaboration and InPTA Collaboration (2023a); Reardon et al. (2023); Xu et al. (2023) — showed that this stochastic process exhibits Hellings-Downs correlations Hellings and Downs (1983) consistent with the isotropic unpolarised stochastic gravitational wave background.
@@
Supermassive black hole binaries at subparsec separations are expected to be a dominant source of the stochastic gravitational wave background at nanohertz frequencies Rosado et al. (2015). However, the expected amplitude of the background is lower than the observations suggest. Previous EPTA analyses found the best-fit strain amplitude to be at the edge of the values simulated from supermassive black hole binary population synthesis models. This is visible in Figure 7 from Nanograv Collaboration (2023b) and Figure 5 from EPTA Collaboration and InPTA Collaboration (2024). The inferred strain amplitude lies at the theoretical upper limit of the predicted astrophysical range Rosado et al. (2015); Casey-Clyde et al. (2022); Simon (2023). It might also be in tension with the observed black hole mass function Sato-Polito et al. (2023); Sato-Polito and Zaldarriaga (2024), although see Liepold and Ma (2024) for a different view. Furthermore, the strain spectral index of the gravitational wave background is in $\approx 2\sigma$ tension with the value corresponding to binary inspirals driven by gravitational wave emission alone. This is visible in Figure 5 in  EPTA Collaboration and InPTA Collaboration (2024) and Figure 11 in  Nanograv Collaboration (2023a). Although previous PTA results were consistent with a very broad range of assumptions about supermassive binary black hole populations Middleton et al. (2018), they suggested deviations from purely gravitational wave-driven binary evolution.
@@
It was pointed out in several studies Goncharov et al. (2021, 2022); van Haasteren (2024a) that the standard PTA models of how noise parameters are distributed across pulsars are incorrect. These models manifest as prior probabilities in PTA data analysis. To be precise, the models are incorrect because they are ‘static’, i.e., the shape of the distribution of noise parameters is not influenced by the data. Although imposing such priors is very unlikely to influence our conclusions about evidence for the gravitational background Zic et al. (2022); Cornish and Sampson (2016); Taylor et al. (2017); Di Marco et al. (2023), it may introduce systematic errors in our measurement of the strain spectrum van Haasteren (2024a).
We address the aforementioned problem. We employ hierarchical Bayesian inference to account for the uncertainties in our noise priors. In particular, we parametrise noise prior distributions which are our models of how parameters governing pulsar-specific noise are distributed across the pulsars van Haasteren (2024a). These newly-introduced hierarchical parameters are called hyperparameters. We use a new procedure of marginalisation over hyperparameters, which is described in Section 2.2.2 of the companion paper Goncharov and Sardana (2024). Furthermore, we discuss other sources of noise that may impact data analysis. Based on the revised noise models, we revisit the measurement of properties of the gravitational wave background with the European Pulsar Timing Array (EPTA) (EPTA Collaboration, 2023a).
[…]
Modelling of pulsar noise is important because it can affect the conclusions about the properties of the gravitational wave background such as $A$ and $\gamma$. Pulsar noise introduces temporal correlations which impossible to distinguish from the effect of gravitational background based solely on single pulsar data. However, unlike for the gravitational background, noise-induced timing delays have different Fourier spectra across pulsars Goncharov et al. (2021). Our hierarchical model describes the ensemble properties of these noise spectra.
Despite similar delay time series induced by both pulsar noise and gravitational background in pulsar data, the presence of the background can still be established. The Hellings-Downs function Hellings and Downs (1983) determines the mean inter-pulsar correlations of the stochastic timing delays induced by the isotropic stochastic unpolarised gravitational background. The statistical significance of the background is established based on the Bayesian odds between (1) the Hellings-Downs-correlated stochastic process with the same spectrum of delays across pulsars and (2) the same spectrum of delays across pulsars and no inter-pulsar correlations Romano et al. (2021).
@@
Results.—Posterior distributions for $A$ and $\gamma$ are shown as contours in Figure 8. Solid blue contours correspond to our improved model. For comparison, dashed red contours correspond to the standard ‘static’ noise priors used in earlier EPTA analyses EPTA Collaboration and InPTA Collaboration (2023a). The results are shown for both the 10-year subset of the EPTA data which showed evidence for the Hellings-Downs correlations, and the full 25-year EPTA data where the evidence is not visible [^2] The correlations are thought to be ‘scrambled’ by unmodelled noise from the older backend-receiver systems of the telescopes.. The value of $\gamma=13/3$ ($\alpha=2/3$) is shown as a dashed straight line. Our improved model results in a lower median-aposteriori strain amplitude of the gravitational wave background, as well as in a steeper spectral index, as shown with solid blue contours in Figure 8. Maximum-aposteriori $(\lg A,\gamma)$ obtained with our improved model and the respective measurement uncertainties based on $1\sigma$ credible levels are reported at the top of Figure 8.
@@
In addition to revising noise priors, we consider the presence of epoch-correlated white noise NANOGrav Collaboration (2016) in the data. Epoch-correlated noise has not been considered in previous EPTA analyses because of a low number (1-4) of pulse arrival time measurements per observation epochs compared to other PTAs EPTA Collaboration and InPTA Collaboration (2023b, a). Nevertheless, we find evidence for this noise in the data from certain backend-received systems. The posterior for our improved model in Figure 1(d) is marginalised over epoch-correlated noise. The choice leads to the consistency of the fully marginalised posterior on $\gamma$ with $13/3$ at $1\sigma$ level. Revising noise priors alone, without considering epoch-correlated noise, yields a weaker consistency of the joint posterior on $(\lg{A},\gamma)$ with $\gamma=13/3$ at $1\sigma$ level (not shown in Figure 1(d)).
(a) $(\lg A,\gamma)=(-14.57^{+0.17}_{-0.34},4.28^{+0.60}_{-0.30})$
@@
Figure 1: Posterior distributions for the power-law amplitude $A$ and spectral index $\gamma$ of the putative gravitational wave background in the European Pulsar Timing Array (EPTA) data. The results of a fit to Hellings-Downs correlations are shown in the left panel (full 25-year data) and the middle panel (the 10-year subset of the data described in ref. EPTA Collaboration and InPTA Collaboration (2023a)). The results of a fit of only temporal correlations to the 10-year data are shown in the right panel. Dashed red contours correspond to the result using the standard pulsar noise priors [^6] Even with the standard noise priors, we report a steeper strain spectral index than in the original EPTA publication EPTA Collaboration and InPTA Collaboration (2023a) because we follow the recommendation by van Haasteren to perform noise model averaging instead of model selection van Haasteren (2024b). , and the blue contours correspond to our improved model [^7] An additional modelling of the epoch-correlated noise is performed for the calculation shown with blue contours in Figure 1(d). [^8] In marginalised distributions, shaded areas correspond to $1\sigma$ credible levels. In joint distributions, an inner dark area corresponds to the $1\sigma$ level, and the outer lighter area corresponds to the $2\sigma$ level.. The horizontal dashed line corresponds to the background from supermassive binary black holes inspiralling entirely due to gravitational wave emission (subject to cosmic variance). Our improved model results in a lower median-aposteriori strain amplitude of the background and mitigates tensions with $\gamma=13/3$.
Implications.—If the energy loss in binary inspirals is dominated by the adiabatic emission of gravitational waves and if binaries are circular, the characteristic strain spectrum of the gravitational wave background is Renzini et al. (2022)
[…]
@@
Tensions of $\gamma\approx 3$ inferred from the previous EPTA analysis with $13/3$ reported during the announcement of evidence for the gravitational background Nanograv Collaboration (2023a); EPTA Collaboration and InPTA Collaboration (2023a); Reardon et al. (2023); Xu et al. (2023) has led to discussions on whether the signal is influenced by certain effects of binary evolution that make strain spectrum to appear flatter. Mechanisms of flattening $h_{\text{c}}(f)$ typically involve the introduction of a more rapid physical mechanism of a reduction in binary separation compared to a gravitational wave emission at $<0.1$ parsec separations. Such a mechanism could be an environmental effect Saeedzadeh et al. (2024) such as stellar scattering Sesana (2010); Sesana and Khan (2015), the torques of a circumbinary gas disc Bonetti et al. (2024). It could also be due to the abundance of binaries in eccentric orbits which lead to a more prominent gravitational wave emission Bi et al. (2023) [^9] Eccentricity also results in a steeper $h_{\text{c}}(f>10^{-8}~{}\text{Hz})$ Burke-Spolaor et al. (2019), but PTA sensitivity declines towards high frequencies.. In contrast, the results of our improved analysis are consistent with binary evolution driven only by the emission of gravitational waves.
@@
Figure 2: The bottom panel shows the predicted log-10 strain amplitude $\lg A$ from $24$ studies. The top panel shows posteriors on $\lg A$ marginalised over $\gamma$. Dashed posteriors are obtained with the original pulsar noise model. Solid posteriors with a band of $1\sigma$ credibility correspond to our improved model [^11] As in Figure 1(d), additional modelling of the epoch-correlated noise is performed for a solid green contour.. The green colour corresponds to the 10-year subset of the EPTA data, the grey colour corresponds to the full 25-year data.
@@
Our improved model also impacts the measurement of the strain amplitude which is directly related to the number density of supermassive black hole binaries. The new results suggest that the supermassive black hole binaries are not as (over-)abundant as the earlier measurements implied. This is shown in Figure 11, which is a replica of Figure A1 in Nanograv Collaboration (2023b). There, green horizontal bands correspond to theoretical uncertainties on the strain amplitude at the $16$th - $84$th percentile level in $24$ studies (Rajagopal and Romani, 1995; Jaffe and Backer, 2003; Wyithe and Loeb, 2003; Enoki et al., 2004; Sesana et al., 2008, 2009; Sesana, 2013; McWilliams et al., 2014; Ravi et al., 2014; Kulier et al., 2015; Ravi et al., 2015; Rosado et al., 2015; Roebber et al., 2016; Sesana et al., 2016; Rasskazov and Merritt, 2017; Dvorkin and Barausse, 2017; Kelley et al., 2017; Ryu et al., 2018; Bonetti et al., 2018; Zhu et al., 2019; Chen et al., 2019, 2020; Siwek et al., 2020; Simon, 2023). Our improved model reduces tensions of the gravitational wave background strain amplitude with theoretical and observationally-based predictions for supermassive black hole binaries. Furthermore, the revised background strain amplitude corresponds to longer delay times between galaxy mergers and supermassive binary black hole mergers Barausse et al. (2023). The caveat is that the reported amplitude is referenced to $f=\text{yr}^{-1}$, so the covariance between $\lg A$ and $\gamma$ in posterior changes following a rotation of a power law about this frequency. Therefore, visible consistency with theoretical predictions may improve based on a different reference frequency.

### v3

# Reading signatures of supermassive binary black holes in pulsar timing array observations

@@ Abstract
We find the inferred properties of the putative gravitational wave background in the second data release of the European Pulsar Timing Array to be in better agreement with theoretical expectations under the improved noise model. In particular, our improved noise models show consistency of the background’s strain spectral index with the value of 2/3-2/3, favoring the population of supermassive black hole binaries as the origin of the background. Our results further suggest that the observed gravitational wave emission is the dominant source of the binary energy loss, with no evidence of environmental effects or eccentric orbits. At the reference gravitational wave frequency of $\text{yr}^{-1}$, we also find a lower power-law strain amplitude of the background than in previous data analyses. This mitigates some of the tensions of the strain amplitude with the expected number density and mass scale of binaries discussed in the literature. However, we show that it is mostly affected by strong covariance of the amplitude and the strain spectral index at $\text{yr}^{-1}$, whereas the strain amplitude at $0.1~\text{yr}^{-1}$ and the strain amplitude at $\text{yr}^{-1}$ assuming a fixed spectral index of 2/3-2/3 remains unaffected. Our results highlight the importance of accurate noise models for correctly inferring properties of the gravitational wave background.
I Introduction
@@
Since 2020, Pulsar Timing Arrays (PTAs) have reported growing evidence for the nanohertz-frequency gravitational wave background in their data. The first tentative evidence came from a temporally-correlated stochastic process in pulsar timing data Nanograv Collaboration (2020); Goncharov et al. (2021a); Chen et al. (2021); Goncharov et al. (2022). The Fourier spectrum of delays and advances in pulsar pulse arrival times exhibited the expected spectral properties of the background. Most recently, PTAs — with varying levels of statistical significance Nanograv Collaboration (2023a); EPTA Collaboration and InPTA Collaboration (2023a); Reardon et al. (2023); Xu et al. (2023) — showed that this stochastic process exhibits Hellings-Downs correlations Hellings and Downs (1983) consistent with the isotropic unpolarised stochastic gravitational wave background.
Properties of the gravitational wave background are assessed using the power law model of its characteristic strain spectrum: $h_{\text{c}}(f)=A(f~\text{yr}^{-1})^{-\alpha}$. Here, $A$ is the strain amplitude at the gravitational wave frequency $f$ of $\text{yr}^{-1}$, and $\alpha$ is the power law spectral index [^1] One may also find a notation where $\alpha^{\prime}=2/3$ and $\gamma=3-2\alpha^{\prime}$.. The corresponding spectral index of the power spectral density $\text{s}^{3}]$ of delays(-advances) $\text{s}]$ induced by the background in the timing data is $\gamma$. These stochastic timing delays resulting from the gravitational-wave redshift of pulsar spin frequency are referred to as temporal correlations.
@@
Supermassive black hole binaries at subparsec separations are expected to be a dominant source of the stochastic gravitational wave background at nanohertz frequencies Rosado et al. (2015). However, the expected amplitude of the background is lower than the observations suggest Mingarelli et al. (2025). Previous European Pulsar Timing Array (EPTA) analyses found the best-fit strain amplitude to be at the edge of the values simulated from supermassive black hole binary population synthesis models. This is visible in Figure 7 from Nanograv Collaboration (2023b) and Figure 5 from EPTA Collaboration and InPTA Collaboration (2024). The inferred strain amplitude lies at the theoretical upper limit of the predicted astrophysical range Rosado et al. (2015); Casey-Clyde et al. (2022); Simon (2023). It might also be in tension with the observed black hole mass function Sato-Polito et al. (2024); Sato-Polito and Zaldarriaga (2025), although see Liepold and Ma (2024) for a different view. Furthermore, the strain spectral index of the gravitational wave background is in $\approx 2\sigma$ tension with the value corresponding to binary inspirals driven by gravitational wave emission alone ($\gamma=13/3$, $\alpha=2/3$) Renzini et al. (2022). The tension is visible in Figure 5 in  EPTA Collaboration and InPTA Collaboration (2024) and Figure 11 in  Nanograv Collaboration (2023a), where the posterior on the spectral index is compared not with a single value, but with the distribution of values expected from realistic backgrounds made up of discrete point sources. Although previous PTA results were consistent with a very broad range of assumptions about binary black hole populations Middleton et al. (2018), they suggested deviations from purely gravitational wave-driven binary evolution. Furthermore, deviations of the strain spectral index from 2/3-2/3 can point to the early-Universe origin of the signal EPTA Collaboration and InPTA Collaboration (2024); Afzal et al. (2023).
@@
Modelling of pulsar-intrinsic noise is important because it can affect the conclusions about the properties of the gravitational wave background, such as $A$ and $\gamma$ (Di Marco et al., 2025). However, despite dedicated noise modelling studies being performed EPTA Collaboration and InPTA Collaboration (2023b), there are four indications that some noise is still mismodelled. First, it was pointed out in several studies Goncharov et al. (2021a, 2022); Goncharov and Sardana (2025) that the standard PTA models of how noise parameters are distributed across pulsars are incorrect. These models manifest as prior probabilities in PTA data analyses. To be precise, the models are incorrect because they are ‘static’, i.e., the shape of the distribution of noise parameters is not influenced by the data. Although imposing such priors is very unlikely to influence our conclusions about evidence for the gravitational background Zic et al. (2022); Cornish and Sampson (2016); Taylor et al. (2017); Di Marco et al. (2023), it may introduce systematic errors in the inferred strain spectrum van Haasteren (2024). Second, epoch-correlated temporally-uncorrelated (white) noise term NANOGrav Collaboration (2016) was neglected in the previous EPTA analysis. Third, van Haasteren (2025) pointed out that the procedure of removing certain noise terms in the search for gravitational waves based on the results of single-pulsar noise analysis, performed in the original EPTA analysis (EPTA Collaboration and InPTA Collaboration, 2023a), is prone to systematic errors. Fourth, the original EPTA analysis treats the transient noise effect in pulsar PSR J1713+0747 as from a sudden change in dispersion measure, although Goncharov et al. (2021b) suggested that it is associated with the pulse shape change.
In this work, we address the above four limitations and revise the properties of the gravitational wave background inferred by the European Pulsar Timing Array (EPTA) (EPTA Collaboration and InPTA Collaboration, 2023a). Based on the revised amplitude and spectral index of the background, we discuss implications for supermassive black hole binaries.
II Results
@@
Posterior distributions for $A$ and $\gamma$ are shown as contours in Figure 3. Solid blue contours correspond to our improved model. For comparison, dashed red contours correspond to the noise model used in the original analysis of the EPTA data EPTA Collaboration and InPTA Collaboration (2023a). The results are shown for both the 10-year subset of the EPTA data which showed evidence for the Hellings-Downs correlations, and the full 25-year EPTA data, where the evidence is not visible (Ferranti et al., 2025a). The value of $\gamma=13/3$ ($\alpha=2/3$) is shown as a dashed straight line. Our improved model results in a lower median-aposteriori strain amplitude of the gravitational wave background, as well as in a steeper spectral index, as shown with solid blue contours in Figure 3. Maximum-aposteriori $(\lg A,\gamma)$ obtained with our improved model and the respective measurement uncertainties based on $1\sigma$ credible levels are reported at the top of Figure 3.
@@
A detailed inspection of the contribution of four new components of our noise models, not shown in Figure 3, reveals the following. Using noise model averaging instead of model selection, as recommended by van Haasteren (2025), has mostly affected the measurement uncertainty. Every other component of our improved noise model results in a consistently lower GWB strain amplitude and a steeper spectral index. Resolving noise prior misspecification using hierarchical inference (Goncharov and Sardana, 2025), the first component of our improved noise model, has mostly affected the 25-year data (Figure 1(b)) and the 10-year data when inter-pulsar correlations are not modelled (Figure 1(f)). When modelling Hellings-Downs correlations in the 10-year data (Figure 1(d)), other components of our revised noise model have made most of the impact. In particular, we found that the amplitude of the transient noise event in PSR J1713+0747 depends on the radio frequency consistently with Goncharov et al. (2021b), not as assumed in the original EPTA analysis EPTA Collaboration and InPTA Collaboration (2023a). Modelling a transient noise event in PSR J1713+0747 correctly results in a significant shift in the posterior. Finally, epoch-correlated noise has not been considered in previous EPTA analyses because of a low number (1-4) of pulse arrival time measurements per observation epoch compared to other PTAs EPTA Collaboration and InPTA Collaboration (2023b, a). Nevertheless, we find evidence for this noise in the data from certain backend-received systems. In Figure 1(d), the choice leads to the consistency of the fully marginalised posterior on $\gamma$ with $13/3$ at $1\sigma$ level. Revising noise priors alone, without considering epoch-correlated noise, yields a weaker consistency of the joint posterior on $(\lg{A},\gamma)$ with $\gamma=13/3$ at $1\sigma$ level (not shown in Figure 1(d)). Stronger impact of hierarchical inference in Figures 1(b) and 1(f) compared to Figure 1(d) illustrates how inter-pulsar correlations influence the posterior.
(a) $(\lg A,\gamma)=(-14.53^{+0.15}_{-0.27},4.14^{+0.56}_{-0.26})$
@@
Figure 1: Posterior distributions for the power-law amplitude $A$ and spectral index $\gamma$ of the putative gravitational wave background in the European Pulsar Timing Array (EPTA) data. The results of a fit to Hellings-Downs correlations are shown in the left panel (full 25-year data) and the middle panel (the 10-year subset of the data described in ref. EPTA Collaboration and InPTA Collaboration (2023a)). The results of a fit of only temporal correlations to the 10-year data are shown in the right panel. Dashed red contours correspond to the result using the standard pulsar noise priors, and the blue contours correspond to our improved model [^3] In marginalised distributions, shaded areas correspond to $1\sigma$ credible levels. In joint distributions, an inner dark area corresponds to the $1\sigma$ level, and the outer lighter area corresponds to the $2\sigma$ level.. The horizontal dashed line corresponds to the background from supermassive binary black holes inspiralling entirely due to gravitational wave emission (subject to cosmic variance). Our improved model results in a lower median-aposteriori strain amplitude of the background and mitigates tensions with $\gamma=13/3$.
II.1 Implications for supermassive binary black holes
If the energy loss in binary inspirals is dominated by the adiabatic emission of gravitational waves and if binaries are circular, the characteristic strain spectrum of the gravitational wave background is Renzini et al. (2022)
$h_{\text{c}}^{2}(f)=\frac{4G^{5/3}}{3\pi^{1/3}c^{2}}f^{-4/3}\int\frac{d^{2}N}{dVdz}\frac{\mathcal{M}^{5/3}}{(1+z)^{1/3}}dz,$ (1)
@@ II.1 Implications for supermassive binary black holes
Tensions of $\gamma\approx 3$ inferred from the previous EPTA analysis with $13/3=4.3(3)$ reported during the announcement of evidence for the gravitational background Nanograv Collaboration (2023a); EPTA Collaboration and InPTA Collaboration (2023a); Reardon et al. (2023); Xu et al. (2023) has led to discussions on whether the signal is influenced by certain effects of binary evolution that make strain spectrum to appear flatter. Mechanisms of flattening $h_{\text{c}}(f)$ typically involve the introduction of a more rapid physical mechanism of a reduction in binary separation compared to a gravitational wave emission at <0.1<0.1 parsec separations. Such a mechanism could be an environmental effect Saeedzadeh et al. (2024) such as stellar scattering Sesana (2010); Sesana and Khan (2015), the torques of a circumbinary gas disc Bonetti et al. (2024). It could also be due to the abundance of binaries in eccentric orbits which lead to a more prominent gravitational wave emission Bi et al. (2023) [^4] Eccentricity also results in a steeper $h_{\text{c}}(f>10^{-8}~\text{Hz})$ Burke-Spolaor et al. (2019), but PTA sensitivity declines towards high frequencies.. In contrast, the results of our improved analysis are consistent with binary evolution driven only by the emission of gravitational waves.
@@
Figure 2: Consistency of the inferred strain amplitude of the gravitational wave background with theoretical expectations. Bottom panels show the predicted log-10 strain amplitude $\lg A$ from $26$ studies. Top panels show posteriors on $\lg A$ marginalised over $\gamma$. Red lines correspond to the original EPTA noise model, blue lines correspond to our revised model. Solid lines correspond to the 10-year data, dashed lines correspond to the 25-year data. Blue bands correspond to $1\sigma$ credible levels for 10-year data and our improved noise model. The left panel shows the characteristic strain amplitude at $\text{yr}^{-1}$, $\lg A$. The right panel shows the strain amplitude at the frequency of $(10~\text{yr})^{-1}$, $\lg A_{10\text{yr}}$.
@@
Our improved model also impacts the measurement of the strain amplitude, which is directly related to the number density of supermassive black hole binaries. The new results suggest that the supermassive black hole binaries are not as (over-)abundant as the earlier measurements implied. This is shown in Figure 2, the bottom panel of which is based on Figure A1 in Nanograv Collaboration (2023b). Grey horizontal bands correspond to theoretical uncertainties on the strain amplitude at the $16$th - $84$th percentile level in $26$ studies (Rajagopal and Romani, 1995; Jaffe and Backer, 2003; Wyithe and Loeb, 2003; Enoki et al., 2004; Sesana et al., 2008, 2009; Sesana, 2013; McWilliams et al., 2014; Barausse, 2012; Barausse et al., 2023; Kulier et al., 2015; Ravi et al., 2015; Rosado et al., 2015; Roebber et al., 2016; Sesana et al., 2016; Rasskazov and Merritt, 2017; Dvorkin and Barausse, 2017; Kelley et al., 2017; Ryu et al., 2018; Bonetti et al., 2018; Zhu et al., 2019; Chen et al., 2019, 2020; Siwek et al., 2020; Simon, 2023). For Barausse (2012) the model is “HS-nod”, and for Barausse et al. (2023) the model is “HS_nod_SN_high_accr”. Our improved model reduces tensions of the gravitational wave background strain amplitude with theoretical and observationally-based predictions for supermassive black hole binaries. Furthermore, the revised background strain amplitude corresponds to longer delay times between galaxy mergers and supermassive binary black hole mergers Barausse et al. (2023). The caveat is that the reported amplitude is referenced to $f=\text{yr}^{-1}$, so the covariance between $\lg A$ and $\gamma$ in posterior changes following a rotation of a power law about this frequency. We tested that at $f=10~\text{yr}^{-1}$ our improved model mostly affects the posterior on $\gamma$, introducing consistency with $13/3$, leaving the posterior on $\lg A_{10\text{yr}}$ almost unchanged. Precisely, with 10-year EPTA data, we measure $(\lg A_{10\text{yr}},\gamma)$ to be $(-14.02^{+0.08}_{-0.11},4.35^{+0.42}_{-1.02})$ with our new noise model and $(-14.00^{+0.08}_{-0.13},3.02^{+0.89}_{-0.25})$ with the original model. Consistently, one may notice in Figure A1 in Nanograv Collaboration (2023b) that the posterior on $\lg A_{10\text{yr}}$ referenced to $f=10~\text{yr}^{-1}$ are in less tension with theoretical predictions.
As shown in Ref. Sato-Polito et al. (2024), Equation 1 can be presented as a function of black hole number density $\rho_{\text{BH}}$, black hole mass scale $M_{*}$, and the intrinsic scatter $\epsilon_{0}$ in galaxy stellar velocity dispersion which is a proxy of mass for supermassive black holes,
$h_{\text{c}}^{2}(f)=1.2\times 10^{-30}\bigg{(}\frac{f}{\text{yr}^{-1}}\bigg{)}^{-4/3}\times\bigg{(}\frac{M_{*}}{5.8\times 10^{7}M_{\odot}}\bigg{)}^{2/3}\times\bigg{(}\frac{\rho_{\text{BH}}}{4.5\times 10^{5}M_{\odot}\text{Mpc}^{-3}}\bigg{)}\times\bigg{(}\frac{2}{e^{\frac{8}{9}\epsilon_{0}^{2}\ln^{2}{(10)}}}\bigg{)}.$ (2)
The mass scale $M_{*}$ is expressed as $\lg M_{*}=a_{\bullet}+b_{\bullet}\frac{\sigma_{*}}{200\text{km~s}^{-1}}$. It is also worth noting that $\rho_{\text{BH}}\propto M_{*}$. Based on the equation above, we perform parameter estimation on $(\rho_{\text{BH}},a_{\bullet},b_{\bullet},\sigma_{*})$ with the 10-year EPTA data and Hellings-Downs correlations. We assume Normal priors for $(a_{\bullet},b_{\bullet},\sigma_{*})$ as well as the value of $\epsilon_{0}=0.38$ from Ref. Sato-Polito et al. (2024). The priors for $(a_{\bullet},b_{\bullet})$ for black hole masses are inferred from kinematic observations of local black holes in galaxy catalogs McConnell and Ma (2013). The prior for $\sigma_{*}$ is based on the velocity dispersion measured in the Sloan Digital Sky Survey Bernardi et al. (2010). We adopt a Normal prior on $\rho_{\text{BH}}$ from Ref. Liepold and Ma (2024) because a prior from Ref. Sato-Polito et al. (2024) yields the strain amplitude in significant tension with the EPTA observations. The posterior and the prior are shown in Figure 3. All parameters are degenerate with each other due to representing a single observed quantity, the strain amplitude. Nevertheless, the figure illustrates which properties of the supermassive black hole population are constrained the most by the pulsar timing data, given astrophysical uncertainties. Provided the scaling in Equation 2, $(0.6\%,5.7\%,0.9\%)$ uncertainties on $(a_{\bullet},b_{\bullet},\sigma_{*})$, which contribute to $M_{*}$, and a larger $44.4\%$ uncertainty on $\rho_{\text{BH}}$, the posterior informs solely on the number density $\rho_{\text{BH}}$. Our improved model influences the posterior to a lesser extent compared to Figure 3 because the spectral index is fixed at $\gamma=13/3$. Accordingly, it can be noticed in Figure 3 that $\lg A$ does not change much for our improved noise model at the slices of fixed $\gamma=13/3$.
Figure 3: Posterior of the population parameters for supermassive black hole binaries with the 10-year EPTA data. Effectively, the constraints are provided solely by the inferred characteristic strain amplitude of the gravitational wave background, assuming the strain power law index of 2/3-2/3. Hellings-Downs correlations are included in the model.
III Discussion